Two different fruit juices are mixed together. 7 liters of the first fruit juice contains 60% apple juice, and 5 liters of the second fruit juice contains 30% apple juice. Find the percentage of the new mix that is apple juice.


1
Expert's answer
2021-08-16T09:41:07-0400

"V_1=7 \;l \\

C_1 = 60 \; \%"

Pure apple juice:

"p_1 = 7 \times 0.60 = 4.2 \;l \\

V_2 = 5\; l \\

C_2 = 30 \; \% \\

p_1 = 5 \times 0.30 = 1.5 \;l \\

V_{total} = 7+5= 12 \;l \\

p_{total} = 4.2+1.5=5.7 \;l"

Proportion:

12 l – 100 %

5.7 l – x

"x=\frac{5.7 \times 100}{12}=47.5 \;%" %

Answer: 47.5 %